OpenStudy (anonymous):

Find the value of the slope of the tangent line to the curve x^2+2xy+3y^2=4 at the point (2,3). Using the Product Rule

OpenStudy (perl):

did you need help with the derivative part?

OpenStudy (anonymous):

I need help with all of it lol I'd like a solution and then an explanation as to how you got there

OpenStudy (perl):

x^2+2xy+3y^2=4 2x + 2( 1*y + x* y ' ) + 3* 2y * y ' = 0

OpenStudy (anonymous):

would it not be \[2x + (2x)(1. \frac{ dy }{ dx }) + (y)(2) + 6y (\frac{ dy }{ dx }) ? \]

OpenStudy (anonymous):

and the 4 becomes a 0

OpenStudy (anonymous):

???

OpenStudy (perl):

i factored out the 2 before doing product rule on x*y

OpenStudy (perl):

2 ( x * y )

OpenStudy (perl):

now use the product rule on x * y

OpenStudy (anonymous):

I don't know how

OpenStudy (anonymous):

show me and then explain

OpenStudy (perl):

(u * v) ' = u ' * v + u * v '

OpenStudy (anonymous):

sorry I don't know how to do it like that. we write it like \[u \frac{ dv }{ dx } + v \frac{ du }{ dx }\]

OpenStudy (anonymous):

Could you do it like that? The other way makes no sense to me

OpenStudy (anonymous):

Sorry this is really important. I need help